2.11 The human eye. The human eye is sensitive from roughly 10-6 cd/m2 (dark night, rod-dominated vision, essentially monochromatic) to about 106 cd/m2 (bright sunlight). Calculate the dynamic range of the eye.
Answer #1

Dynamic range of the eye is ratio of brightest luminance to darkest luminance. Dynamic range = (10^6)/ (10^-6)= 10^12(10 power 12).